# Alt-MoE:Multimodal Alignment via Alternating Optimization of Multi-directional MoE with Unimodal Models

### v1

@@ Abstract
Recent Large Multi-Modal Models (LMMs) have made significant advancements in multi-modal alignment by employing lightweight connection modules to facilitate the representation and fusion of knowledge from existing pre-trained uni-modal models. However, these methods still rely on modality-specific and direction-specific connectors, leading to compartmentalized knowledge representations and reduced computational efficiency, which limits the model’s ability to form unified multi-modal representations. To address these issues, we introduce a novel training framework, Alt-MoE, which employs the Mixture of Experts (MoE) as a unified multi-directional connector across modalities, and employs a multi-step sequential alternating unidirectional alignment strategy, which converges to bidirectional alignment over iterations. The extensive empirical studies revealed the following key points: 1) Alt-MoE achieves competitive results by integrating diverse knowledge representations from uni-modal models. This approach seamlessly fuses the specialized expertise of existing high-performance uni-modal models, effectively synthesizing their domain-specific knowledge into a cohesive multi-modal representation. 2) Alt-MoE efficiently scales to new tasks and modalities without altering its model architecture or training strategy. Furthermore, Alt-MoE operates in latent space, supporting vector pre-storage and real-time retrieval via lightweight multi-directional MoE, thereby facilitating massive data processing. Our methodology has been validated on several well-performing uni-modal models (LLAMA3, Qwen2, and DINOv2). achieving competitive results on a wide range of downstream tasks and datasets.
@@ 1 Introduction
Figure 1: Paradigm of our Alt-MoE joint embedding predictive architecture. The self-supervised learning is conducted with two encoding branches of input and output signals, and a multi-directional MoE predictor to match the input latent embedding into the target latent embedding by minimizing contrastive and prediction losses. The MoE predictor is conditioned on the multi-modal routed information which is learned with information entropy minimization. Training is performed with alternative any-to-any multi-modality tasks.
Human perception is inherently multi-modal, seamlessly integrating diverse sensory inputs from vision, hearing, touch, and other senses to comprehend the world. Inspired by this capability, multi-modal learning aims to develop Artificial Intelligence (AI) systems that can process and interpret multiple types of input simultaneously, thereby mimicking human-like cognition. multi-modal large-scale models can process and integrate information from multiple modalities such as text, images, audio and video, and have become an important way to solve complex tasks involving heterogeneous data sources and achieve general artificial intelligence Alayrac et al. (2022); Radford et al. (2021); Wang et al. (2022a, 2023); Gao et al. (2024).
Modern multi-modal studies are mostly built on self-supervised learning (SSL) and generative architecture, while can be classified into two main categories. The first category aim to learn the multi-modal information by training from scratch Wang et al. (2022a, 2023); Oquab et al. (2023). Such methodology asks for large-scale data therefore faces high training costs. It may also limit the acquisition of intra-modal information and introduces inter-modality information bias. For example, vision-language pre-training studies might be predominantly text-guided Radford et al. (2021), which hinders the retainment of visual information due to inherent limitations of text descriptions Oquab et al. (2023). Another category either employs the pretrained Large Language Model (LLMs) as the backbone Wu et al. (2023); Haotian Liu (2023); Jun Zhan (2024) or a lightweight cross-modal connector Alayrac et al. (2022); Junnan Li (2023); Bai et al. (2023) in order to preserve fine-grained uni-modality information and alleviate computational burdens Zhang et al. (2024). Success of such methodologies indicates the "platonic representation" Huh et al. (2024), which indicates representations across various domains are increasingly converging to the same latent space. Such representation similarity spans different model architectures, training objectives, and even data modalities.
Typical generative architecture tries to reconstruct the target signal conditioned by the input signal directly, usually with a decoder structure Assran et al. (2023). However, such approach is not good at eliminating irrelevant details, and may face difficulty when input details are not easily predictable, or representing complex uncertainties in continuous spaces Dawid and LeCun (2023). It will also create bias given ambiguous semantic alignment of multi-modal samples. To take a step ahead on the representation convergence on multi-modalities, one may consider to reconstruct multi-modal signals in the latent-space, however, which might induce the representation collapse, especially when the latent variables contain excess information Dawid and LeCun (2023). Such consideration calls for a different framework with better alignment in the latent space, and reduced information content of latent variables.
In this work, we propose a new multi-modal paradigm which studies the dependency of the unobserved part ($y$) on the observed part ($x$) in their embedding space, in which $x$ and $y$ may belong to different modalities or any combination of them. We employ the joint-embedding predictive architecture (JEPA) Dawid and LeCun (2023); Assran et al. (2023) to align the representation of different modalities into the same space. JEPA achieves the alignment by first encoding the observed and unobserved parts, then use a predictor to convert one embedding into another embedding space, driven by the contrastive and predictive losses simultaneously. We implement the predictor by recap the lightweight cross-modal connector, which is designed as the Mixture-of-Experts (MoE) structure. We solve the framework by Alternating Gradient Descent (AGD) of different multi-modal tasks, similar with IMP Akbari et al. (2023). To tackle the challenge of the multi-modal semantic discrepancy (paired cross-modal data may exhibit inconsistent information despite their intended correspondence), we decouple the cross-modal information into modality-independent and shared components through distinct experts. This multi-directional MoE is used to conduct multiple unidirectional alignment tasks and is solved by Alternating Gradient Descent (AGD) of different multi-modal tasks, similar with IMP Akbari et al. (2023). The router information of MoE is the latent factor $l$ which helps better alignment. To avoid the representation collapse, we solve the problem by minimizing the information content of predictor Dawid and LeCun (2023). Our Alt-MoE paradigm is depicted by Figure 1.
@@
To empirically validate these properties and explicitly evaluate the alignment effectiveness, we conducted comprehensive experiments across multiple multi-modal tasks, while keeping the model architecture and training strategy consistent. Experimental results demonstrate that our Alt-MoE easily achieves modality and task scaling while attaining competitive results compared to current state-of-the-art (SoTA) multi-modal studies. Furthermore, our approach is able to conduct large-scale on-line multi-modal retrieval tasks due to our unique architecture, while mainstream multi-modal models fail to do so. We summarize our contributions as follows:
@@
    We propose Alt-MoE, a novel modality-agnostic multi-modal alignment method that leverages multi-directional MoE as a cross-modal connector. Through alternating gradient descent, Alt-MoE alternates between each unidirectional alignment task and optimizing the multi-directional MoE, enabling efficient and scalable alignment across arbitrary modalities.
@@
    Alt-MoE achieves multi-modal alignment in the latent space, which avoids the modality information bias and facilitates large-scale online retrieval ability.
  3. 3.
    Unlike conventional direct alignment methods, Alt-MoE decomposes complex multi-modal tasks into tractable unidirectional sub-tasks, employing alternating optimization for global convergence. This approach offers a new theoretical framework for multi-modal research, potentially enhancing efficiency and scalability across diverse modalities.
  4. 4.
    We provide an information-theoretic explanation in the problem formulation, with extensive experimental results and theoretical analysis supporting its effectiveness in improving alignment and handling complex multi-modal interactions.
  5. 5.
    Our experimental results demonstrate that effective multi-modal alignment can be achieved by using lightweight connectors to link uni-modal models. Furthermore, Alt-MoE operates in latent space, which enables vector pre-storage and real-time retrieval, thereby facilitating efficient processing of large-scale data tasks.
The rest of the paper is organized as follows. The connection with previous works is first discussed in Section 2. Problem is formulated in Section 3. The methodology is stated in Section 4. Experiment results are summarized in Section 5. We also discuss current limitations and future plans in Section 6. Finally Section 7 concludes this paper. Project code and model checkpoints can be found in https://github.com/HongyangLL/Alt-MoE.
@@ 2.1 End-to-end multi-modal Learning
Recently, end-to-end multi-modal models employing various architectures have achieved outstanding performance. These architectures can be broadly categorized into several main types: Dual-encoder Radford et al. (2021); Jia et al. (2021). Fusion-encoder Li et al. (2020); Jia et al. (2021); Chen et al. (2020). Combining dual encoder and fusion encoder architectures integrate specialized layers into multi-modal models to enable deep cross-modal interactions Li et al. (2021, 2022).
The majority of multi-modal learning method employ large-scale multi-modal datasets for end-to-end pre-training. However, as model scale continues to increase, several potential challenges emerge: Firstly, the pre-training process may incur prohibitively high computational costs. Secondly, these models often struggle to adapt to novel modalities or tasks without extensive retraining. Moreover, most multi-modal approaches employ text-guided learning of visual concepts, which can limit the acquisition of fine-grained information inherent to each modality. A logical approach is to leverage existing pre-trained uni-modal foundation models Li et al. (2023); Oquab et al. (2023); Zhang et al. (2024).
Unlike these works, each step in Alt-MoE close to I-JEPA Assran et al. (2023). I-JEPA learn to predict the embeddings of a signal $y$ from a compatible signal $x$, using a predictor network that is conditioned on additional (possibly latent) variables $z$ to facilitate prediction. However, due to the complexity of multi-modal alignment, relying solely on a predictive structure proves insufficient for effective alignment. To enhance cross-modal retrieval capabilities, Alt-MoE incorporates a contrastive structure into Alt-MoE framework. Furthermore, to achieve alternating alignment and decouple cross-modal information, we replace the predictor with a multi-directional MoE. Cross embeddings serve as the signal $z$, prompting the MoE to execute tasks across different modalities.
@@ 2.2 Multi-modal Learning with uni-modal Models
Recent trends in multi-modal learning have increasingly focused on integrating high-performance uni-modal models to achieve effective multi-modal capabilities. Flamingo Alayrac et al. (2022) integrates visual information into each layer of a frozen Large LLM through the use of cross-attention. BLIP-2 Li et al. (2023) introduces an additional vision-to-language adaptation module, Q-former, and proposes a two-stage training process to mitigate the challenges associated with learning vision-language alignment. However, both methods require substantial parameters and multi-modal data for cross-modal alignment.
Recent studies indicate a convergence of representations across modalities, providing evidence for the feasibility of developing advanced multi-modal models by connecting high-performance uni-modal models with lightweight parameters and data-efficient techniques Huh et al. (2024). For instance, LLaVA Liu et al. (2024) achieved state-of-the-art performance by employing a two-layer multilayer perceptron (MLP). Similar architectures have subsequently proliferated across various domains Li et al. (2024); Zhang et al. (2024); Lin et al. (2024, 2023). Alt-MoE further advances this concept by interconnecting diverse high-performance uni-modal models through a shared multi-directional MoE. We conducted extensive experiments focused on alignment efficacy, demonstrating the effectiveness of joint training across multiple modalities and directions.
@@ 2.3 Multi-modal learning with MoE and AGD
Prior studies have investigated AGD-based multi-modal multi-task alternating training, revealing that the integration of diverse modalities, tasks, and resolutions can yield mutual benefits, thereby effectively enhancing the model’s generalization capabilities and cross-domain performance Akbari et al. (2023); Likhosherstov et al. (2021). we further extend this approach to integrate existing pre-trained high-performance uni-modal models, achieving overall alignment through alternating bidirectional alignment.
MoE-LLaVA Lin et al. (2024) proposes MoE-Tuning, a strategy for Large Vision-Language Models that creates a sparse model with constant computational cost. While both MoE-LLaVA and Alt-MoE employ sparse MoE to connect high-performance uni-modal large models, Alt-MoE not only scales this approach to large audio-visual-language models but also achieves pairwise bidirectional alignment across modalities.
@@ 3 Problem Formulation
In this section, we will first introduce the overall optimization objective of multi-modal bidirectional alignment, followed by the optimization objective of unidirectional alignment, and finally the alignment loss and theoretical explanation. It is worth noting that by integrating uni-modal models, Alt-MoE can effectively align multiple modalities. To clearly articulate the design rationale, we illustrate this with the example of image-text alignment.
@@
    Multi-modal alignment objective: Alt-MoE decomposes multi-modal alignment into multiple unidirectional alignment subtasks, where MoE serves as a knowledge representation and fusion module and achieve multi-modal alignment by alternately aligning from image to text and text to image.
@@
    We provide an information-theoretic explanation to elucidate the rationale behind the design of the training objective.
Next, we will describe the optimization process at two levels, including the optimization objectives and parameter updates.
@@ 3.2 Alternating Unidirectional Alignment
In this section, we delineate the optimization objectives and parameter update procedures across various time steps $t$. By introducing AGD, we can alternately optimize unidirectional alignment at each time step $t$ with the goals of maximizing mutual information between image $I$ and text $T$, denoted as $I(I;T)$, and minimizing conditional entropies, denoted as $H(T|I)$ and $H(I|T)$.
@@ 3.3 Information Decomposition and Alignment
Figure 2: Overview of Alt-MoE: Alt-MoE connects different modalities through multi-directional MoE and performs unidirectional alignment from image to text and text to image at different time steps, ultimately achieving overall bi-directional alignment. The specific structure of the multi-directional MoE is shown in the figure on the right. Different embeddings prompt the MoE to select different experts to perform contrastive learning or prediction.
@@ 4 Methodology
In this section, we will provide a detailed introduction to the architecture of Alt-MoE. As shown in Figure 2, Alt-MoE is divided into three modules: the visual model (VM), the language model (LM), and the fusion module MoE. We input the paired images and text $(I,T)$ into VM and LM respectively to obtain latent representations $z_{i},z_{t}$, and then perform multi-modal interaction in the fusion module.

### v3

# Alt-MoE:A Scalable Framework for Bidirectional Multimodal Alignment and Efficient Knowledge Integration

@@ Abstract
Multimodal learning has advanced significantly by aligning different modalities within shared latent spaces, enabling tasks such as cross modal understanding and generation. Current alignment strategies in multimodal learning primarily include direct alignment using pre-trained or unified encoders and single-directional alignment via modality-specific connectors. Direct alignment struggles to fully leverage rich intra-modal knowledge, often requiring extensive training data to achieve cross-modal representation. Meanwhile, single-directional alignment methods, despite leveraging pre-trained knowledge, restrict task adaptability and hinder the model’s ability to capture bidirectional relationships, leading to incomplete knowledge fusion and underutilization of complementary modality-specific information. To address these limitations, we introduce Alt-MoE, a scalable multimodal alignment framework that employs a mixture of experts (MoE) model as a multi-directional connector across modalities. By utilizing a sequential alternating one-way alignment strategy, Alt-MoE iteratively refines the model to achieve bidirectional alignment. Alt-MoE operates in latent space, enabling efficient vector pre-storage and real-time retrieval via MoE, optimizing large-scale data processing. Extensive empirical studies demonstrate that Alt-MoE achieves competitive performance on cross-modal retrieval and visual question answering by integrating diverse modality-specific knowledge, generalizing to unseen data, and easily scaling to new tasks and modalities through dynamic adjustment of MoE capacity and expert activation.
@@ 1 Introduction
Figure 1: Primary strategies for aligning modalities: Figure (a). Direct Alignment Framework: This framework illustrates two types of alignment. On the left, a dual-encoder approach is shown, where an x-encoder and a y-encoder independently process inputs $x$ and $y$, producing representations $s_{x}$ and $s_{y}$. A similarity function $D(s_{x},s_{y})$ is used to align the two modalities directly. On the right, a unified multimodal encoder jointly processes $x$ and $y$, producing a fused representation $s_{xy}$. The training objective combines two loss functions: $L_{itm}$, focusing on image-text matching, and $L_{lm}$, which corresponds to the language modeling loss. Figure (b). Unidirectional Alignment Framework: In this framework, an x-encoder processes input $x$ to produce a representation $s_{x}$, which is then refined by a predictor. $I$ provides additional information to guide the predictor in aligning refined representations with a large language model (LLM). The LLM is trained with the language modeling loss $L_{lm}$, using input $y$ as the target. This framework focuses on aligning the modality $x$ unidirectionally to the language modality through the LLM. Figure (c.) Alt-MoE models cross-modal bidirectional alignment by alternately executing unidirectional alignment tasks in the latent space (detail in 4.3). $s_{x}$, $s_{y}$ represent the encoding results of the $x$-encoder and $y$-encoder, respectively. $t$, $t+1$ represent the time steps. $L$ denotes the loss function (prediction and contrastive loss). $z_{x2y}$ and $z_{y2x}$ represent the cross embeddings (detail in 4.3).
Achieving human-level intelligence, a central goal of Artificial General Intelligence (AGI), requires building hierarchical world models capable of abstract reasoning. These models enable agents to predict, plan, and reason about complex, multi-layered environments. Abstract reasoning, particularly in high-dimensional latent spaces, is crucial for cross-modal understanding and generation. It requires integrating diverse sources of information across multiple modalities to construct unified, high-level representations. These representations must transcend individual modalities, enabling models to capture complex, abstract relationships [22, 8].
However, current multimodal learning often fall short in effectively capturing and utilizing this abstract reasoning across modalities. Current multimodal learning frameworks often rely on two primary strategies for aligning modalities: direct alignment using pre-trained or unified multimodal encoders and single-directional alignment through modality-specific connectors. Direct alignment methods, such as CLIP [37] and ImageBind [13], involve aligning modalities by mapping them into a shared space using pre-trained encoders. VLMo [4] and ONE-PEACE [40] utilize a unified multimodal encoder to integrate and model interactions across different modalities. On the other hand, single-directional alignment strategies, like BLIP-2 [25] and LLaVA [33], focus on connecting specific modalities through specialized components or connectors, often operating in a one-way alignment direction. These approaches aim to establish coherent cross-modal representations, but each comes with its limitations in capturing the full complexity of multimodal relationships.
Direct alignment of multiple encoders excel in generalization, and efficient cross-modal retrieval [25, 42, 13, 47]. Unified multimodal encoder facilitates deeper interactions across modalities and is particularly advantageous in scenarios requiring joint reasoning and holistic understanding [41, 40, 4]. Nevertheless, both approaches share limitations: they often miss fine-grained intra-modal details, struggle with tasks that demand abstract reasoning, and rely heavily on large-scale pre-training data. In text-guided visual representation learning, captions offer an abstract view of images, often missing pixel-level details crucial for complex visual reasoning. This limitation challenges direct alignment methods in retaining fine-grained intra-modal features, restricting their deep visual understanding capabilities [35].
Another method employs single-directional alignment through modality-specific connectors, such as LLaVA [33] and BLIP-2 [26]. These methods connect different modalities via specialized components, aligning them in a single direction. While this approach facilitates efficient cross-modal interaction and preserves unimodal richness, it can limit the model’s ability to capture complex, bidirectional relationships necessary for abstract reasoning. As a result, while it excels in specific tasks, it may struggle with more intricate, abstract reasoning that requires deeper intermodal dependencies.
Building upon these observations, we introduce Alt-MoE, a novel multi-modal alignment framework. Alt-MoE employs a unified, multi-directional MoE as cross-modal connector, enabling any-to-any modality alignment. This framework integrate the rich intra-modal knowledge of pre-trained models, preventing the loss of fine-grained information while avoiding high computational costs and data requirements. By employing Alternating Gradient Descent (AGD), bidirectional alignment decomposes into alternating unidirectional tasks, ultimately achieving bidirectional alignment. This method not only mitigates the risks of incomplete information and abstract reasoning deficiencies inherent in single-directional alignment but also allows the model to incorporate multiple loss functions, achieving more robust cross-modal alignment. Alt-MoE introduces a paradigm shift by exploring latent dependencies, enabling more dynamic and abstract forms of reasoning.
Further expanding on this, Alt-MoE is a new multi-modal paradigm which studies the dependency of the unobserved part ($y$) on the observed part ($x$) in their latent space. This enables the model to explore different hypothetical scenarios, plan actions, or make inferences about unobserved parts of the world based on observed data, thus facilitating more sophisticated and generalized reasoning. Experts within the MoE can operate at different levels of abstraction, ensuring that each modality is processed in a way that suits its nature while maintaining coherence in the overall model.
@@
To assess the model’s performance, we achieved superior results across multiple tasks (e.g., VQA, image-text retrieval, audio-text retrieval, image classification) and datasets. Additionally, we validated the model’s generalization on unseen data by testing its performance on zero-shot audio-text retrieval. Experimental results demonstrate that our Alt-MoE easily achieves modality and task scaling while attaining competitive results compared to current state-of-the-art (SoTA) multi-modal studies. Furthermore, our approach is able to conduct large-scale on-line multi-modal retrieval tasks due to our unique architecture. We summarize our contributions as follows:
@@
    Alt-MoE introduces a modality-agnostic alignment method that reduces training and data costs while leveraging the rich knowledge in pre-trained models. This approach offers a new framework for multi-modal research, potentially enhancing efficiency and scalability across diverse modalities.
@@
    Alt-MoE leverages multi-directional MoE for cross-modal alignment in latent space, building a hierarchical world model that abstracts complex dependencies.
@@ 2.1 End-to-end Multi-modal Learning
Recently, end-to-end multi-modal models employing various architectures have achieved outstanding performance. These architectures can be broadly categorized into several main types: Dual-encoder [37, 18]. unified-encoder [27, 18, 6]. Combining dual encoder and fusion encoder architectures integrate specialized layers into multi-modal models to enable deep cross-modal interactions [24, 25].
The majority of multi-modal learning method employ large-scale multi-modal datasets for end-to-end pre-training. However, as model scale continues to increase, several potential challenges emerge:d Firstly, the pre-training process may incur prohibitively high computational costs. Secondly, these models often struggle to adapt to novel modalities or tasks without extensive retraining. Moreover, most multi-modal approaches employ text-guided learning of visual concepts, which can limit the acquisition of fine-grained information inherent to each modality. A logical approach is to leverage existing pre-trained uni-modal foundation models [26, 35, 45].
@@ 2.2 Multi-modal Learning with Uni-modal Models
Recent trends in multi-modal learning have increasingly focused on integrating high-performance uni-modal models to achieve multi-modal alignment. Flamingo [2] integrates visual information into each layer of a frozen Large LLM through the use of cross-attention. BLIP-2 [26] introduces an additional vision-to-language adaptation module, Q-former, and proposes a two-stage training process to mitigate the challenges associated with learning vision-language alignment. However, both methods require substantial parameters and multi-modal data for cross-modal alignment.
Recent studies indicate a convergence of representations across modalities, providing evidence for the feasibility of developing advanced multi-modal models by connecting high-performance uni-modal models with lightweight parameters and data-efficient techniques [16]. For instance, LLaVA [33] achieved state-of-the-art performance by employing a two-layer multilayer perceptron (MLP). Similar architectures have subsequently proliferated across various domains [23, 45, 30, 29]. Alt-MoE further advances this concept by interconnecting diverse high-performance uni-modal models through a shared multi-directional MoE. We conducted extensive experiments focused on alignment efficacy, demonstrating the effectiveness of joint training across multiple modalities and directions.
@@ 2.3 Multi-modal Learning with MoE and AGD
Prior studies have investigated AGD-based multi-modal multi-task alternating training, revealing that the integration of diverse modalities, tasks, and resolutions can yield mutual benefits, thereby effectively enhancing the model’s generalization capabilities and cross-domain performance [1, 28]. we further extend this approach to integrate existing pre-trained high-performance uni-modal models, achieving overall alignment through alternating bidirectional alignment.
MoE-LLaVA [30] proposes MoE-Tuning, a strategy for Large Vision-Language Models that creates a sparse model with constant computational cost. While both MoE-LLaVA and Alt-MoE employ sparse MoE to connect high-performance uni-modal large models, Alt-MoE not only scales this approach to large audio-visual-language models but also achieves pairwise bidirectional alignment across modalities.
@@ 3 Problem Formulation
In this section, we first introduce the overall optimization objective of bidirectional alignment, followed by the optimization objective of unidirectional alignment, and finally the alignment loss and theoretical explanation. To clearly articulate the design rationale, we illustrate this with the example of image-text alignment.
@@
    Multi-modal alignment objective: Alt-MoE decomposes multi-modal bidirectional alignment into multiple unidirectional alignment subtasks, where MoE serves as a cross modal knowledge representation and fusion module and achieve multi-modal alignment by alternately aligning from image to text and text to image.
@@
    Information Decomposition and Alignment: we provide an information-theoretic explanation to elucidate the rationale behind the design of the training objective.
@@ 3.2 Alternating Unidirectional Alignment
In this section, we delineate the optimization objectives and parameter update procedures across various time steps $t$. By introducing AGD, we can alternately optimize unidirectional alignment at each time step $t$ with the goals of maximizing mutual information between image $I$ and text $T$, denoted as $I(I;T)$, and minimizing conditional entropies, denoted as $H(T|I)$ and $H(I|T)$. The details of the multi-directional MoE are provided in the supplementary materials.
@@ 3.3 Information Decomposition and Alignment
Figure 2: Overview of Alt-MoE: Alt-MoE connects different modalities through multi-directional MoE and performs unidirectional alignment from image to text and text to image at different time steps, ultimately achieving overall bi-directional alignment. Here, we take image-text alignment as an example. The left part of the figure illustrates the process of image-text alignment, where we alternately perform image-to-text and text-to-image alignment on a unified multi-directional MoE. The specific structure of the multi-directional MoE is shown in the figure on the right. Different embeddings prompt the MoE to select different experts to perform contrastive learning or prediction. CL refers to Contrastive Learning, and Dual Pred denotes Bidirectional Prediction. To guide the MoE in performing different sub-tasks, we introduce modality embeddings (e.g., text embeddings, image embeddings). These embeddings are incorporated through cross-embedding mechanisms to effectively steer the tasks.
@@ 4 Methodology
In this section, we take image-text alignment as an example to provide a detailed introduction to the architecture of Alt-MoE. As shown in Figure 2, Alt-MoE is divided into three modules: the visual model (VM), the language model (LM), and the bidirectional alignment module MoE. We input the paired images and text $(I,T)$ into VM and LM respectively to obtain latent representations $z_{i},z_{t}$, and then perform multi-modal interaction in the fusion module.
4.1 Image and Text Encoding
$z_{i}$ $=f_{v}(I),\quad z_{i}\in\mathbb{R}^{i},$ (6)
$z_{t}$ $=f_{l}(T),\quad z_{t}\in\mathbb{R}^{t},$